# Rendezvous in CAVITY: Kinematics and gas properties of an isolated dwarf-dwarf merging pair in a cosmic void region

### v1

@@ Appendix A More on the kinematics of the system
As discussed in Section 4.1, a lopsided rotation is observed in both the gas and stellar radial velocity maps of this merging system. Part of this lopsidedness can be attributed to the 100 km s-1 difference in the relative velocity between the merging components. Additionally, the higher values in a small linear region between the components, as previously mentioned, are likely due to the emission lines from both components overlapping in a single spaxel, which are not resolved by the PPAK IFU’s spectral and spatial resolution. However, this region serves as a useful separation line between the components. In Fig. 9, we show the radial velocity maps of each merging component, separated by the elevated line. After accounting for the radial velocity difference, component B exhibits clear rotation. Component A also shows signs of rotation, although less clearly, possibly due to its alignment along the line of sight and the significant effects of the interaction on its kinematics.
@@
In Table 4, we present the input and output parameters used for the fitting for each merging component, as described in Section 4.2. The details regarding the parameters and their meaning can be found in the documentation page of : https://bbarolo.readthedocs.io/en/latest/tasks/fit3d.html#parameters. Here, we assume that the position of the center (xpos, ypos), the scale-height of the disc (z0), and the systemic velocity (vsys) remain constant across the discs of the galaxies. To simplify the fitting process, we provided initial values for these parameters to , and kept them fixed during the fitting. For vsys, we did not specify a value; instead, it was estimated by the software as the central velocity from the global line profile.
@@
norm | AZIM | AZIM | Default/Recommended value |
@@
redshift | 0 | 0 | Default/Recommended value |
linear | - | - | Default/Recommended value |
bweight | 0 | 0 | Default/Recommended value |
normalcube | FALSE | FALSE | Default/Recommended value |
@@
The columns display the names of both the input and output parameters, along with the corresponding values for components A and B, and relevant comments for each parameter. Parameters that were fixed during the fit are marked with an asterisk and are listed in the first part of the table, alongside other input parameters. The second part of the table presents the geometrical parameters derived from the fit results. The comments are categorized into four general groups: Fine-tuning are parameters estimated after several iterations. Datacube limitations are parameters determined as the best value for the datacube. Isophote fitting are parameters estimated using isophote fitting. Default/Recommended value are parameters set as default or recommended by the pipeline.

### v2

@@ Appendix A More on the kinematics of the system
As discussed in Section 4.1, a lopsided rotation is observed in both the gas and stellar radial velocity maps of this merging system. Part of this lopsidedness can be attributed to the 100 km s-1 difference in the relative velocity between the merging components. Additionally, the higher values in a small linear region between the components, as previously mentioned, are likely due to the emission lines from both components overlapping in a single spaxel, that are not resolved by the PPAK IFU’s spectral and spatial resolution. However, this region serves as a useful separation line between the components. In Fig. 9, we show the radial velocity maps of each merging component, separated by the elevated line. After accounting for the radial velocity difference, component B exhibits clear rotation. Component A also shows signs of rotation, although less clearly, possibly due to its alignment along the line of sight and the significant effects of the interaction on its kinematics.
@@
In Table 4, we present the input and output parameters used for the fitting for each merging component, as described in Section 4.3. The details regarding the parameters and their meaning can be found in the documentation page of 666https://bbarolo.readthedocs.io/en/latest/tasks/fit3d.html#parameters. Here, we assumed that the position of the centre(xpos, ypos), the scale-height of the disc (z0), and the systemic velocity (vsys) remain constant across the discs of the galaxies. To simplify the fitting process, we provided initial values for these parameters to , and kept them fixed during the fitting. For vsys, we did not specify a value; instead, it was estimated by the software as the central velocity from the global line profile.
@@
| norm | AZIM | AZIM | Default or recommended value |
@@
| redshift | 0 | 0 | Default or recommended value |
| linear | - | - | Default or recommended value |
| bweight | 0 | 0 | Default or recommended value |
| normalcube | FALSE | FALSE | Default or recommended value |
@@
Input and output parameters of . The columns display the names of both the input and output parameters, along with the corresponding values for components A and B, and relevant comments for each parameter. Parameters that were fixed during the fit are marked with an asterisk and are listed in the first part of the table, alongside other input parameters. The second part of the table presents the geometrical parameters derived from the fit results. The comments are categorised into four general groups: Fine-tuning are parameters estimated after several iterations. Datacube limitations are parameters determined as the best value for the datacube. Isophote fitting are parameters estimated using isophote fitting. Default or recommended value are parameters set as default or recommended by the pipeline.